A coin tossed and die is rolled. The outcome of the coin is recoded "1" when it shows a head, and "0" when it shows a tail. The random variable gives the sum of the outcomes of coin and die. Compute the average value of the random variable?

1
Expert's answer
2021-03-02T05:20:55-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & 1 & 0 \\ \hline
 1 & 1+1=2 & 1+0=1 \\
 \hdashline
2 & 2+1=3 & 2+0=2 \\
 \hdashline
3 & 3+1=4 & 3+0=3 \\
 \hdashline
4 & 4+1=5 & 4+0=4 \\
 \hdashline
5 & 5+1=6 & 5+0=5 \\
 \hdashline
6 & 6+1=7 & 6+0=6 \\
 
\end{array}"

"E(X)=mean=\dfrac{1}{12}(1)+\dfrac{2}{12}(2)+\dfrac{2}{12}(3)"

"+\dfrac{2}{12}(4)+\dfrac{2}{12}(5)+\dfrac{2}{12}(6)+\dfrac{1}{12}(7)=4"

The average value of the random variable is 4.